Alex is 6 years older than Sara. Sixteen years ago, Alex was 1 1/2 times as old as Sara. How old are they now?
Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)


Hi
Alex is 6 years older than Sara
Let x and (x+6) represent the ages of Sara and Alex NOW respectively
Question states*** 16 years ago...
(x+6) - 16 = 3/2(x-16)
solving for x
x - 10 = 1.5x - 24
14 = .5x
14/.5 = x
x = 28, Sara's age NOW. Alex is 34 NOW. (28 + 6)
CHECKING our Answer***
18 = 1.5 * 12 = 18

Answer by stanbon(75887)

Alex is 6 years older than Sara. Sixteen years ago, Alex was 1 1/2 times as old as Sara. How old are they now?
---------
Equations:
a = s+6
(a-16) = (3/2)(s-16)
-----
Substitute for "a" and solve for "s":
(s+6-16) = (3/2)(s-16)
2(s-10) = 3s-48
2s-20 = 3s-48
s = 28 (Sara's age now)
----------------------
Solve for "a":
a = s+6
a = 28+6
a = 34 (Alex's age now)
============================
cheers,
Stan H.


--------------------------------------------------------------------------------